1. Are the following lines parallel, perpendicular, or neither?
L1 with equation x – 2y = 10
L2 with equation 2x + y = 2

2. Given f(x) = –x3 – 3x2 + 3x + 7, find f(–1), f(0), and f(2)
Many thanks in advance since I am doing my finals!

Answer by stanbon(75887)

1. Are the following lines parallel, perpendicular, or neither?
L1 with equation x – 2y = 10
Slope-intercept form:
L1: y=(x-10)/2, so slope is 1/2
L2 with equation 2x + y = 2
L2: y=-2x+2, so slope is -2
Since the slopes are the negative-reciprocal of each other
the lines are perpendicular.
2. Given f(x) = –x^3 – 3x^2 + 3x + 7, find f(–1), f(0), and f(2)
Substitute to find f(x):
f(-1)=-(-1)^3-2(-1)^2+3(-1)+7
=1-2-3+7=3
Do the same for the other parts of this problem.
Cheers,
Stan H.